OpenStudy (anonymous):

how many cups is 1080 ml in fraction form?

OpenStudy (anonymous):

Depends on the Cup size, if it's 250ml, then it's \[\frac{ 1080 }{ 250 } ,\] But the aim of reducing fractions down further, is to get them to the smallest fraction possible by using numbers that both the top/bottom or numerator/denominator will divide by evenly, with no decimals. The easiest one's are /2 /3 /4, Let's test this here. \[\frac{ 1080 }{ 250 } /2 = \frac{ 540 }{ 125 }\] \[\frac{ 1080 }{ 250 } /3 = \frac{ 360 }{ 83.3 }\] ^ Decimals so we won't use this. \[\frac{ 1080 }{ 250 } /4 = \frac{ 270 }{ 62.5 }\] ^ Decimal so we won't use this. Let's go back to and see if we can make that even smaller. \[\frac{ 1080 }{ 250 } /2 = \frac{ 540 }{ 125 }\] \[\frac{ 540 }{ 125 }/5=\frac{ 108 }{ 25 }\]As you can see, it doesn't look like this can be reduced any further, so your final answer is 108/25.

OpenStudy (shane_b):

For a standard U.S. "cup" of liquid it's going to be\[1080\cancel{mL}*\frac{1cup}{236.588\cancel{mL}}=?\space cups\]

OpenStudy (shane_b):

Obviously you can reduce the final answer there.

OpenStudy (anonymous):

Another way to think about it is:\[\frac{ X }{ Y }/A\] X is your initial value the one you want to divide into the Y value, in this case a container. Then you dived X/A and Y/A, Your A value is whatever is the largest number that X/Y can divide into evenly. (No decimals) In your example earlier X= 1080 and Y = 250, A was 2 ( At first ) then A was 5. I like to reduce sometimes by an initial number. If you notice, \[\frac{ 1080 }{ 250 } /10 = \frac{ 108 }{ 25 },\]Which was the same result as dividing it by 2 then by 5, and as you already have probably noticed 2x5=10, the value that 1080/250 can divide into the most evenly and reduced down to the most.